# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

### v2

@@ 4 Sublinear convergence of stochastic PDR for restricted primal-dual gap
$=\mathbb{E}[\sum_{n=0}^{k}p_{i}(1-p_{i})^{n}\varphi(\tilde{s}_{i}^{k+1-n})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$ (4.4)
which leads to (i). For (ii), the trick is to find the relations among those three items. That is, for every $k$, $\|\tilde{s}^{k+1}-s^{k}\|^{2}=\|s^{k+1}-s^{k}\|^{2}+\|\tilde{s}^{k+1}-s^{k+1}\|^{2}$ since $\|s_{i}^{k+1}-s_{i}^{k}\|^{2}$ will be zero except $i=i_{k}$ and $\|\tilde{s}_{i}^{k+1}-s_{i}^{k+1}\|^{2}$ will be zero except $i\neq i_{k}$. Using $\mathbb{E}\|s^{k+1}-s^{k}\|^{2}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k}\|_{P}^{2}$, the statement (ii) follows. Now, let us prove $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test},i}-y_{i}^{*}\|^{2}$ is uniformly bounded for all $k$.
$\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test},i}-y_{i}^{*}\|^{2}$ $=\mathbb{E}\|\tilde{\bar{y}}_{i}^{k+1}-\bar{y}_{i}^{k}+\tilde{y}_{i}^{k+1}-y_{i}^{*}\|^{2}\leq\mathbb{E}\|\tilde{\bar{y}}^{k+1}-\bar{y}^{*}+\bar{y}^{*}-\bar{y}^{k}+\tilde{y}^{k+1}-y^{*}\|^{2}$
$\leq 3\mathbb{E}\|\tilde{\bar{y}}^{k+1}-\bar{y}^{*}\|^{2}+3\mathbb{E}\|\tilde{y}^{k+1}-y^{*}\|^{2}+3\tau\mathbb{E}\|\bar{y}^{k}-\bar{y}^{*}\|_{\tau^{-1}P^{-1}}^{2}$
$\leq 3\mathbb{E}\|\mathcal{T}u^{k}-u^{*}\|^{2}+3\tau\mathbb{E}\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p}}$
$\leq(3C+3\tau)\mathbb{E}\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p}}\leq(3C+3\tau)\Delta^{0}$ (4.5)
where the third inequality is deduced from Lemma 5 while the last one is from (4.1). Applying (i) with $\varphi(\cdot)=(\cdot-y^{*}_{i})^{2}$ on $y^{k+1}_{\operatorname{test},i}$ and noting $10<\sum_{n=0}^{k}p_{i}(1-p_{i})^{n}<$ 1, $(1-p_{i})^{k+1}<1$, we arrive at
$\mathbb{E}\|y^{k+1}_{\operatorname{test},i}-y_{i}^{*}\|^{2}$ $=\mathbb{E}\sum_{n=0}^{k}{p_{i}(1-p_{i})^{n}}\|\tilde{y}^{k+1-n}_{\operatorname{test},i}-y_{i}^{*}\|^{2}+(1-p_{i})^{k+1}\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2},$
$\leq(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}.$ (4.6)
Summing (4.6) over all the components for $i=1,\ldots,n$ leads to (iii). Now, let us turn to (iv). Remembering the updates of $\tilde{y}^{k+1}_{\operatorname{test}}$, $\tilde{y}^{k+1}$, $x^{k+1}$ as in (SPDR) and the non-expansiveness of the resolvent operator $(I+\tau\partial\mathcal{G})^{-1}$, it can be readily checked that
$\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}\leq\mathbb{E}\sum_{k=1}^{K}\|(2\tilde{y}^{k+1}-\bar{y}^{k})-(2\tilde{y}^{k}-\bar{y}^{k-1})\|^{2}$
@@
$D=4\max({\sigma(\|A_{1}\|\|(N_{1}-I)\|)^{2},\ {\sigma}\|A_{1}\|^{2},\ \tau(\|A_{2}\|\|(N_{2}-I)\|)^{2},\ {\tau}\|A_{3}\|^{2})}$ (4.8)
Combining (4.7) and (4.1), we obtain (iv). For (v), since $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|_{P}^{2}$, we henceforth focus on $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}$. By the inequality $\|a+b\|^{2}\leq(1+\epsilon)\|a\|^{2}+(1+{1}/{\epsilon})\|b\|^{2}$, for any $\epsilon>0$ and $K\geq 1$, we arrive at
@@
where the first equality is obtained by (ii). For the convenience of proofs, letting $\epsilon={1}/{\underline{p}}$ with $\underline{p}:=\min_{i}\{p_{i},\ i=1,\ldots,n\}$ henceforth, then $(1+{1}/{\epsilon})(I-\mathbf{P})\leq(1+{1}/{\epsilon})(1-\underline{p})I=(1-\underline{p}^{2})I$. For employing Lemma 6, choosing $\delta=\underline{p}^{2}$ and $S_{0}=0$, we denote for $K\geq 1$,
@@
$\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}\leq\mathbb{E}\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}=\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|^{2}_{(1-\delta)I}.$ (4.11)
@@
$h_{K}\leq(1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}$ (4.12)
and $\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{1-\delta}^{2}$ is given by (4.11). Thus by (4.10), (4.12), Lemma 6 and the fact $S_{0}=0$, we have
$S_{K}$ $\leq\frac{1-(1-\delta)^{K-1}}{\delta}\left((1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)\leq\frac{1}{\delta}\left((1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right).$
@@
$\leq\frac{1-(1-\delta)^{K-1}}{\delta}\left((1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
Finally, with (ii) and $\delta=\underline{p}^{2}$, $\epsilon=1/\underline{p}$, we derive
@@
$\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+1/\underline{p})D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{K},$ (4.13)
$\leq{1}/{\underline{p}^{2}}\left((1+1/\underline{p})D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{\infty}.$
Now, let us extend a lemma in [1] originated from [19] to positive semidefinite weights.
@@
With notations $s^{k}$ defined in Lemma 4, picking up a point $s^{0}_{\dagger}\in\mathbb{Y}$ at will, let us define an iteration sequence $v^{k+1}=s^{k}-\tilde{s}^{k+1}-\mathbf{P}^{-1}(s^{k}-s^{k+1})$ with ${s}_{\dagger}^{k+1}={s}_{\dagger}^{k}-\mathbf{P}v^{k+1}$. Then, it follows that for any $s\in\mathbb{Y}$ and a positive semidefinite, self-adjoint, and linear weight $\mathfrak{D}=\emph{Diag}[\mathfrak{D}_{1},\ldots,\mathfrak{D}_{n}]$ with $\mathfrak{D}_{i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ and $\mathfrak{D}\mathbf{P}=\mathbf{P}\mathfrak{D}$ as before,
$\sum_{k=0}^{K}\langle{s}_{\dagger}^{k}-s,v^{k+1}\rangle_{\mathfrak{D}}=\frac{1}{2}\|{s}_{\dagger}^{0}-s\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{s}_{\dagger}^{K+1}-s\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2},$ (4.14)
and
$\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2}]\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|s^{k}-s^{k+1}\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}].$ (4.15)
@@
$\frac{1}{2}\|{s}_{\dagger}^{k+1}-s\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}$ $=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}-\langle\mathbf{P}v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathbf{P}^{-1}\mathfrak{D}}+\frac{1}{2}\|\mathbf{P}v^{k+1}\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}$
$=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}-\langle v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathfrak{D}}+\frac{1}{2}\|v^{k+1}\|_{\mathbf{P}\mathfrak{D}}^{2}.$ (4.17)
Summing the equality (4), we arrive at (4.14). For the second result, similar to $\mathbb{E}^{k}\|v^{k+1}\|^{2}=\mathbb{E}^{k}\|\mathbb{E}^{k}\xi-\xi\|^{2}\leq\mathbb{E}^{k}\|\xi\|^{2}$, with $\xi^{k}=\mathbf{P}^{-1}(s^{k}-s^{k+1})$ and $\mathbb{E}^{k}\xi=s^{k}-\tilde{s}^{k+1}$, we obtain
$\mathbb{E}[\sum_{k=0}^{K}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2}]=\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2}]]$
$\leq\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|\mathbf{P}^{-1}(s^{k+1}-s^{k})\|_{\mathfrak{D}\mathbf{P}}^{2}]]=\sum_{k=0}^{K}\mathbb{E}[\|s^{k+1}-s^{k}\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}].$ (4.18)
@@
With the preceding lemmas, we now prove the sublinear convergence of (SPDR). The main ideas are employing the relation between the stochastic Bregman distance $\mathcal{H}(z_{\operatorname{test},K},z)$ and the deterministic Bregman distance $\mathcal{H}(\tilde{z}_{\operatorname{test},K},z)$ together with separating the operation $\mathbb{E}$ from the operation $\sup_{z\in\mathbb{B}}$ when acting on $\mathcal{H}(z_{\operatorname{test},K},z)$.
@@
If $u=(x,y,\bar{x},\bar{y}),$ with $\bar{x}=x+\sigma\mathcal{K}^{\ast}y,$ $\bar{y}=y-\tau\mathcal{K}x$, and $z^{k}_{\operatorname{test}}=(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, a bounded domain, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K},$ (4.19)

### v3

@@ 4 Sublinear convergence of stochastic RPDR for restricted primal-dual gap
$=\mathbb{E}[\sum_{j=0}^{k}p_{i}(1-p_{i})^{j}\varphi(\tilde{s}_{i}^{k+1-j})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$ (4.4)
which leads to (i).
For (ii), the trick is to find the relations among those three items. That is, for every $k$, $\|\tilde{s}^{k+1}-s^{k}\|^{2}=\|s^{k+1}-s^{k}\|^{2}+\|\tilde{s}^{k+1}-s^{k+1}\|^{2}$ since $\|s_{i}^{k+1}-s_{i}^{k}\|^{2}$ will be zero except $i=i_{k}$ and $\|\tilde{s}_{i}^{k+1}-s_{i}^{k+1}\|^{2}$ will be zero except $i\neq i_{k}$. Using $\mathbb{E}\|s^{k+1}-s^{k}\|^{2}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k}\|_{P}^{2}$, the statement (ii) follows. Now, let us prove $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ is uniformly bounded for all $k$, based on (3.4) and the iteration formula of PDR, we have
$\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ $=\mathbb{E}\|\tilde{y}^{k+1}_{t,\operatorname{test}}-y^{*}\|^{2}=\mathbb{E}\|\tilde{\bar{y}}_{t}^{k+1}-\bar{y}^{k}+\tilde{y}_{t}^{k+1}-y^{*}\|^{2}$
$\leq 3\mathbb{E}\|\tilde{\bar{y}}_{t}^{k+1}-\bar{y}^{*}\|^{2}+3\mathbb{E}\|\tilde{y}_{t}^{k+1}-y^{*}\|^{2}+3\mathbb{E}\|\bar{y}^{k}-\bar{y}^{*}\|^{2}$
$\leq 3\mathbb{E}\|\mathcal{T}u^{k}-u^{*}\|^{2}+3\tau\rho_{k,y}\mathbb{E}\|\bar{y}^{k}-\bar{y}^{*}\|_{\tau^{-1}\rho_{k,y}^{-1}\mathbf{P}^{-1}}^{2}$
$\leq 3C\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}^{2}+3\tau\rho_{u}\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}$
$\leq 3C\rho_{u}\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}+3\tau\rho_{u}\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}$
$=(3C+3\tau)\rho_{u}\mathbb{E}\Delta_{k}\leq(3C+3\tau)\rho_{u}\Delta^{0}\leq\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}$
where the third inequality is deduced from Lemma 5 while the last one is from (4.1). Applying (i) with $\varphi(\cdot)=\|\cdot-y^{*}\|^{2}$ on $y^{k+1}_{\operatorname{test}}$ and noting $10<\sum_{j=0}^{k}p_{i}(1-p_{i})^{j}<$ 1, 0<($10<(1-p_{i})^{k+1}<$ 1, then $0\prec\sum_{j=0}^{k}\mathbf{P}(I-\mathbf{P})^{j}\prec I$, $0\prec(I-\mathbf{P})^{k+1}\prec I$, we arrive at
$\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ $=\mathbb{E}\sum_{j=0}^{k}\|\tilde{y}^{k+1-j}_{\operatorname{test}}-y^{*}\|_{\mathbf{P}(I-\mathbf{P})^{j}}^{2}+\|y^{0}_{\operatorname{test}}-y^{*}\|_{(I-\mathbf{P})^{k+1}}^{2},$
$\leq\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2}.$ (4.5)
Now, let us turn to (iv). Remembering the relation between $\tilde{y}^{k+1}_{\operatorname{test}}$ and $\tilde{y}^{k+1}_{t,\operatorname{test}}$ as in (3.4), the updates of PDR and the non-expansiveness of the resolvent operator $(I+\tau\partial\mathcal{G})^{-1}$, it can be readily checked that
$\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{t,\operatorname{test}}-\tilde{y}^{k}_{t,\operatorname{test}}\|^{2}\leq\mathbb{E}\sum_{k=1}^{K}\|(2\tilde{y}_{t}^{k+1}-\bar{y}^{k})-(2\tilde{y}_{t}^{k}-\bar{y}^{k-1})\|^{2}$
@@
$D=4\max({\sigma(\|A_{1}\|^{2}\|N_{1}-I\|),\ {\sigma}\|A_{1}\|^{2},\ \tau\|A_{2}\|^{2}\|N_{2}-I\|,\ {\tau}\|A_{3}\|^{2}}).$ (4.7)
Combining (4.6) and (4.1), we obtain (iv).
For (v), since $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|_{P}^{2}$, we henceforth focus on $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}$. By the inequality $\|a+b\|^{2}\leq(1+\epsilon)\|a\|^{2}+(1+{1}/{\epsilon})\|b\|^{2}$, for any $\epsilon>0$ and $K\geq 1$, we arrive at
@@
where the first equality is obtained by (ii). For the convenience of proofs, letting $\epsilon={1}/{\underline{p}}$ with $\underline{p}:=\min_{i}\{p_{i},\ i=1,\ldots,n\}$ henceforth, then $(1+{1}/{\epsilon})(I-\mathbf{P})\leq(1+{1}/{\epsilon})(1-\underline{p})I=(1-\underline{p}^{2})I$.
For employing Lemma 6, choosing $\delta=\underline{p}^{2}$ and $S_{0}=0$, we denote for $K\geq 1$,
@@
$\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}\leq\mathbb{E}\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}=\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|^{2}_{(1-\underline{p}^{2})I}.$ (4.10)
@@
$h_{K}\leq(1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}.$ (4.11)
Thus by (4.9), (4.11), Lemma 6 and the fact $S_{0}=0$, we have
$S_{K}$ $\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)$ (4.12)
$\leq\frac{1}{\underline{p}^{2}}\left((1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right).$
@@
$\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
Finally, with (ii), we derive
@@
$\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+1/\underline{p})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{K},$ (4.13)
$\leq{1}/{\underline{p}^{2}}\left((1+1/\underline{p})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{\infty}.$
Let us extend a lemma in [1] originated from [34] to positive semidefinite weights.
@@
With notations $s^{k}$ defined in Lemma 4, picking up a point $s^{0}_{\dagger}\in\mathbb{Y}$ at will, let us define an iteration sequence $v^{k+1}=s^{k}-\tilde{s}^{k+1}-\mathbf{P}^{-1}(s^{k}-s^{k+1})$ with ${s}_{\dagger}^{k+1}={s}_{\dagger}^{k}-\mathbf{P}v^{k+1}$. Introduce a positive semidefinite, self-adjoint, and linear weight $\mathfrak{D}=\emph{Diag}[\mathfrak{d}_{1},\ldots,\mathfrak{d}_{n}]$ with $\mathfrak{d}_{i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$, and a non-increasing positive weight $\Theta^{k}=\emph{Diag}[\vartheta_{1}^{k}I,\ldots,\vartheta_{n}^{k}I]$ with $\vartheta_{i}^{k}I:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ and $0<\vartheta_{l}\leq\vartheta_{i}^{k+1}\leq\vartheta_{i}^{k}\leq\vartheta_{u}$ for $i=1,\ldots,n$ and $k=0,\ldots,K-1$. Here $\vartheta_{l}$ and $\vartheta_{u}$ are positive constants. We also assume that the weights $\Theta^{k}$, $\mathfrak{D}$ and the probability matrix $\mathbf{P}$ can communicate with each other, i.e., $\mathbf{P}\Theta^{k}=\Theta^{k}\mathbf{P}$, $\mathfrak{D}\Theta^{k}=\Theta^{k}\mathfrak{D}$ together with $\mathfrak{D}\mathbf{P}=\mathbf{P}\mathfrak{D}$ as before. Then, it follows that for any $s\in\mathbb{Y}$
$\sum_{k=0}^{K}\langle{s}_{\dagger}^{k}-s,v^{k+1}\rangle_{\mathfrak{D}\Theta^{k}}\leq\frac{1}{2}\|{s}_{\dagger}^{0}-s\|_{\mathfrak{D}\Theta^{0}\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{s}_{\dagger}^{K+1}-s\|_{\mathfrak{D}\Theta^{K}\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\Theta^{k}\mathbf{P}}^{2},$ (4.14)
where the above inequality in (4.14) becomes equality while $\Theta^{k}\equiv\Theta^{0}$. We further have
$\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\Theta^{k}\mathbf{P}}^{2}]\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|s^{k}-s^{k+1}\|_{\mathfrak{D}\Theta^{k}\mathbf{P}^{-1}}^{2}].$ (4.15)
@@
$\frac{1}{2}\|{s}_{\dagger}^{k+1}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}$ $=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}-\langle\mathbf{P}v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}+\frac{1}{2}\|\mathbf{P}v^{k+1}\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}$
$=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}-\langle v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathfrak{D}\Theta^{k}}+\frac{1}{2}\|v^{k+1}\|_{\mathbf{P}\mathfrak{D}\Theta^{k}}^{2}.$ (4.17)
Summing the equality (4), we arrive at
$\sum_{k=0}^{K}\langle{s}_{\dagger}^{k}-s,v^{k+1}\rangle_{\mathfrak{D}\Theta^{k}}=\frac{1}{2}\sum_{k=0}^{K}\left[\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}-\|{s}_{\dagger}^{k+1}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}\right]+\frac{1}{2}\sum_{k=0}^{K}\|v^{k+1}\|_{\mathbf{P}\mathfrak{D}\Theta^{k}}^{2}.$
Noting that $\Theta^{k}\succeq\Theta^{k+1}$ together with $\mathbf{P}\mathfrak{D}\Theta^{k}\succeq\mathbf{P}\mathfrak{D}\Theta^{k+1}$ for $k=0,\ldots,K-1$, we obtain (4.14) from the above equation. For the second result, similar to $\mathbb{E}^{k}\|v^{k+1}\|^{2}=\mathbb{E}^{k}\|\mathbb{E}^{k}\xi-\xi\|^{2}\leq\mathbb{E}^{k}\|\xi\|^{2}$, with $\xi^{k}=\mathbf{P}^{-1}(s^{k}-s^{k+1})$ and $\mathbb{E}^{k}\xi=s^{k}-\tilde{s}^{k+1}$, we obtain
$\mathbb{E}[\sum_{k=0}^{K}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}\Theta^{k}}^{2}]=\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}\Theta^{k}}^{2}]]$
$\leq\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|\mathbf{P}^{-1}(s^{k+1}-s^{k})\|_{\mathfrak{D}\mathbf{P}\Theta^{k}}^{2}]]=\sum_{k=0}^{K}\mathbb{E}[\|s^{k+1}-s^{k}\|_{\mathfrak{D}\mathbf{P}^{-1}\Theta^{k}}^{2}].$ (4.18)
@@
With the preceding lemmas, we now prove the sublinear convergence of SRPDR.
@@
If $u=(x,y,\bar{x},\bar{y}),$ with $\bar{x}=x+\sigma\mathcal{K}^{\ast}y,$ $\bar{y}=y-\tau\mathcal{K}x$. Let $u^{k}$ be the iteration sequence generated by SRPDR, and let $z^{k}_{\operatorname{test}}=(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ be the corresponding transitional variables. Denote $K_{1}:=K+1$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K+1}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}=\frac{1}{K_{1}}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, a bounded domain that contains a saddle-point, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K_{1}}=\mathcal{O}(\frac{1}{K}),$ (4.19)